OpenStudy (anonymous):

x2 + 24x + 144 Part 1: Explain how the polynomial is factored. Part 2: Provide the factors

OpenStudy (anonymous):

Have you heard of the sum and product method? Or something similar in wording?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well ok. It is basically like this: x^2 is the 1st term 24x is the 2nd term 144 is the 3rd term Do you understand this?

OpenStudy (anonymous):

Now you have to look at the 3rd term (144). What two numbers multiply to get the 2nd term (24x)? (just forget about the x for the moment)

OpenStudy (anonymous):

There are many possiblities: 12 * 12 24 * 6 72 * 2 etc Which ones of these multiplication factors ADD to get the 2nd term? eg 12+12=24 24+6= 24 72+2=24

OpenStudy (anonymous):

Then you have the brackets (x + ?) (x +?) When expanding you know that x * x =x^2. Now those two factors that are multiplied to get the 3rd term and are added to get the 2nd term, goes where the ? are.

OpenStudy (anonymous):

so (x+12)(x+12). To double check you got it right, expand: =x^2 +12x +12x +144 =x^2+24x+144 Therefore it is correct.